Determine the spacing distance (in mm) of grating contain 317 slits per centimeter.
Answer #1

As here in the question there is no other information except the slit per centimeter.

The slit per cm , N= 317 per cm = 31.7 per mm

Thus the spacing distance, d is = 1/N = 1/31.7 in mm = 0.0315 mm.